how to integrate from(0 to pi) for x/(1+sin^2x) please help
 

Grade:12

2 Answers

Saurabh Koranglekar
4 years ago
\dfrac{-\ln\left(3-2^\frac{3}{2}\right)\ln\left(\left|\mathrm{e}^{2\mathrm{i}x}+2^\frac{3}{2}-3\right|\right)+\ln\left(2^\frac{3}{2}+3\right)\ln\left(\left|\mathrm{e}^{2\mathrm{i}x}-2^\frac{3}{2}-3\right|\right)+\operatorname{Li}_2\left(1-\left(2^\frac{3}{2}+3\right)\mathrm{e}^{2\mathrm{i}x}\right)-\operatorname{Li}_2\left(\left(2^\frac{3}{2}-3\right)\mathrm{e}^{2\mathrm{i}x}+1\right)}{2^\frac{5}{2}}

Regards
Vikas TU
4 years ago
Dear student 
Ans will be 
I = Pi 
Hope this will help …..................
Good Luck …...........
Cheers